# Nonisotropic Gaussian Diffusion for Realistic 3D Human Motion Prediction

### v1

@@ 4.2 SkeletonDiffusion
SkeletonDiffusion is an LDM composed of a conditional autoencoder and a denoiser network, trained in two different stages (Fig. 3). First, we train the autoencoder to reconstruct future motions Y by encoding them in the latent space $\boldsymbol{z}=\mathrm{{e}}({\textbf{Y}})\in\mathbb{R}^{J\times L}$ and decoding them conditioning on the last two observation frames as $\mathrm{{d}}(\boldsymbol{z},\textbf{X}_{-2:0})$. After, the denoiser network $\mathrm{{g}}$ learns with our nonisotropic diffusion to denoise true latent variables $\boldsymbol{z}$ conditioned on past observations $\mathrm{{e}}(\textbf{X})$. At inference, we recover multiple latent codes ${\boldsymbol{x}}_{\theta}$ by sampling and denoising isotropic Gaussian noise and decode them into future predictions $\tilde{{\textbf{Y}}}$.
Latent Representation
While our nonisotropic diffusion has been defined in Sec. 4.1 to operate on $\boldsymbol{z}\in\mathbb{R}^{J}$, we aim for a richer latent representation, capable of encoding the complex temporal information. Hence, we opt for a two-dimensional latent representation $\boldsymbol{z}\in\mathbb{R}^{J\times L}$ where every $j$-th body joint is described by a vector of dimension $L$. Notice that the latent feature dimension $L$ does not explicitly encode information between joints. Thus we assume i.i.d noise over this dimension and perform nonisotropic diffusion only along the $J$ one. Similar latent configurations have been shown effective in other domains, e.g., in image generation [62], but have not been applied to HMP before. In Sup.Mat., we show the correlation of the latent joint dimensions through PCA.
@@
We advocate for an architecture that performs attention on the skeleton joints [65] from end to end throughout the whole model and opt for Typed-Graph (TG) Convolutions [65]. For a TG layer taking as input features ${\boldsymbol{x}}\in\mathbb{R}^{J\times D_{in}}$, we define a feature extraction matrix $\mathbf{W}^{j}\in\mathbb{R}^{D_{in}\times D_{out}}$ for each joint $j$ with shared weights depending on the specific joint, and a feature aggregation matrix $\mathbf{G}\in\mathbb{R}^{J\times J}$. The features $\mathbf{f}\in\mathbb{R}^{J\times D_{out}}$ are first extracted for each joint $j$ independently through
$\mathbf{f}=\mathbf{G}\cdot\hat{\mathbf{f}},\,\text{with}\,\,\hat{\mathbf{f}}^{j}=\mathbf{W}^{j}\cdot{\boldsymbol{x}}^{j}$ (8)
and then aggregated. We further define the TG version of multi-head self-attention [73] as Typed-Graph Attention (TG-Attention). By obtaining and aggregating the query, key, and value matrices through TG layers, we apply the self-attention mechanism on a joint level (equations in Sup.Mat.). Both encoder and decoder are TG-GRUs [65], exploiting the convenient inductive biases of recurrent neural networks (RNNs) for motion modeling [49]. The denoiser network is designed as a TG-Attention network with residual connections. In this way, we define a fully TG architecture that preserves the body joint dimensions continuously throughout the whole architecture.
Training Objective
First, the conditional autoencoder is trained to encode and reconstruct future motions according to the autoencoder objective:
$L_{\mathrm{autoenc}}=L_{\mathrm{rec}}(\mathrm{{e}}({\textbf{Y}}),{\textbf{Y}},\textbf{X}),$ (9)
@@
$L_{\mathrm{rec}}(\boldsymbol{z},{\textbf{Y}},\textbf{X}):=\|\mathrm{{d}}(\boldsymbol{z},\textbf{X}_{-2:0})-{\textbf{Y}}\|_{1}.$ (10)
Here the conditioning on the past encourages smooth transitions between past and future motions. More details on the first training stage can be found in Sup.Mat. After the first training stage, the learned latent joint dimensions exhibit correlations similar to skeleton body joints - this is the foundation of our nonisotropic diffusion training. In the second training stage, the denoising network $\mathrm{{g}}$ learns to reverse this noisification process. Instead of predicting the noise ${\boldsymbol{\epsilon}}_{t}$ [31, 62], we aim to directly approximate the true latent code ${\boldsymbol{x}}_{0}:=\boldsymbol{z}$ [61, 5] as ${\boldsymbol{x}}_{\boldsymbol{\theta}}=\mathrm{{g}}({\boldsymbol{x}}_{t},\mathrm{{e}}(\textbf{X}),t)$. The diffusion loss can be formulated as a Mahalanobis distance:
$L_{\mathrm{diff}}({\boldsymbol{x}}_{\boldsymbol{\theta}},{\boldsymbol{x}}_{0},t):=\bar{\alpha}_{t}\|\bar{\boldsymbol{\Lambda}}_{t}^{-1/2}\boldsymbol{U}^{\top}({\boldsymbol{x}}_{\boldsymbol{\theta}}-{\boldsymbol{x}}_{0})\|^{2},$ (11)
To implicitly enforce diversity [5, 67, 29], we relax Eq. (11) by sampling $k=50$ predictions at each iteration and backpropagating the gradient only through the prediction that minimizes the loss.
@@
Figure 4: Qualitative Results on AMASS. On top, we report the input past observation and the ground-truth future. For every method, we report the closest prediction to GT according to ADE on the right and the two furthermost predictions on the two left-most columns. Our closest competitors generate realistic motions but do not include a motion close to the GT (BeLFusion) or diverse predictions, and present evident unrealistic artifacts (CoMusion). Our method is the only one to produce realistic and diverse motions.
@@ 5.1 Experimental Settings
We evaluate on the large-scale dataset AMASS [52] according to the cross-dataset evaluation protocol [5, 67]. We aim to test SHMP methods with real-world data obtained not from MoCap but from noise sources (e.g., RGB cameras, and sparse IMUs). To this end, we perform zero-shot experiments on 3D Poses in the Wild (3DPW)[76] for available models trained on AMASS, and adapt the recent in-the-wild, large-scale dataset FreeMan[79] to the motion prediction task and retrain on it various state-of-the-art methods. Compared to AMASS, we deem the conventionally employed Human3.6M dataset [34] less representative (only 7 subjects) and discuss it directly in Sup.Mat. with further details. As in previous works, we predict the next 2s into the future from observations of 0.5s.
@@
Recent SHMP works concentrate on four factors: precision, coverage of the ground truth test distribution (multimodal metrics), diversity, and realism. We consider conventionally employed metrics [88, 5] and report their definition and discussion in Sup.Mat.
It is worth noting that while the CMD metric addresses realism, it is solely expressed in terms of joint velocities. The actual Body realism, e.g., bone lengths preservation along the motion, although crucial for meaningful predictions, is overlooked. Even worse, a consequence of artifacts such as changes in limb lengths over the timespan (limbs stretching) and frequent inconsistencies between consecutive frames (limbs jitter) impact other metrics, for example, by causing more diversity in the predictions, and so higher APD value (see Sup.Mat. for further discussion). This motivates us to investigate this aspect and propose novel metrics. Given a future ground truth sequence Y with $\mathrm{B}$ limbs (or bones) and a predicted sequence $\tilde{{\textbf{Y}}}$, for each frame $\tau$ of the prediction associated pose $\tilde{\textbf{p}}_{\tau}$, we denote the length of the $j$-th limb as $\tilde{b}^{j}_{\tau}\in\mathbb{R}$. We define the normalized $j$-th limb length error $e^{j}_{\tau}$ and limb jitter $v^{j}_{\tau}$ at a time $\tau$ as:
$e^{j}_{\tau}:=\frac{1}{b^{j}}\left|b^{j}-\tilde{b}^{j}_{\tau}\right|,\quad v^{j}_{\tau}:=\frac{1}{b^{j}}\left|\tilde{b}^{j}_{\tau+1}-\tilde{b}^{j}_{\tau}\right|$ (13)
where $b^{j}\in\mathbb{R}$ is the ground truth length of the $j$-th limb. By calculating the mean and root mean square error (RMSE) of these two vectors over the time dimension, we define four body realism metrics: mean for stretching str and jitter jit, and analogously RMSE.
@@
HumanMAC [16] 0.415 0.511 0.537 0.600 5.426 2.025 7.91 1.49 11.89 1.84
BeLFusion [5] 0.420 0.495 0.496 0.516 5.209 6.306 10.46 0.41 11.93 0.54
CoMusion [67] 0.389 0.480 0.527 0.525 6.687 2.764 7.94 0.81 10.27 1.053
SkeletonDiffusion (Ours) 0.374 0.457 0.506 0.508 6.732 3.166 7.58 0.51 9.64 0.66
@@ 5.2 Results
Following the cross-dataset evaluation protocol of BelFusion [5], we train on a subset of datasets belonging to AMASS and test on others. We report our results in Tab. 1. Starting from the conventional metrics evaluation, our method already achieves state-of-the-art performance on the majority of the metrics, with a significant improvement on precision. Among other Diffusion-based methods (DM), SkeletonDiffusion and CoMusion contend with each other for first and second place according to diversity, realism, and multimodal metrics. Although VAE-methods tend to show higher diversity, as already mentioned by previous works [5, 16], the high diversity values of [82, 88] may often be the consequence of unrealistic motions with irregularities between past and future or inconsistent speed. From the qualitative example reported in Fig. 4, we notice that both the most diverse predictions of DivSamp represent a cartwheeling motion. While such motions may geometrically be diverse from each other and thus increase the APD, they are not only not semantically diverse but also not consistent with the past observation.
@@
Figure 7: Motions’ validity on different error tolerance on AMASS. For every method, we show the evolution of valid motions quantity (y-axis) for which the maximal error is below a given threshold (x-axis). SkeletonDiffusion presents consistently the highest number of valid poses. CoMusion and VAE methods cannot generate predictions with an error lower than 2.5%.
@@
VAE TPK [77] 0.648 0.701 0.665 0.702 9.582 13.136 7.61 0.36 10.02 0.51
DLow [88] 0.581 0.649 0.602 0.651 13.772 11.977 8.53 0.42 11.28 0.61
GSPS [54] 0.552 0.650 0.578 0.653 11.809 12.722 6.38 0.29 8.65 0.35
DivSampling [19] 0.554 0.678 0.593 0.686 24.153 46.431 11.04 0.78 16.31 1.01
DM BeLFusion [5] 0.493 0.590 0.531 0.599 7.740 17.725 6.47 0.22 7.96 0.29
CoMusion 0.477 0.570 0.540 0.587 11.404 7.093 4.01 0.38 5.54 0.50
DM SkeletonDiffusion (Ours) 0.472 0.575 0.535 0.594 9.814 10.474 3.02 0.17 4.16 0.23
@@
ADE $\downarrow$ FDE $\downarrow$ MMADE $\downarrow$ MMFDE $\downarrow$ APD $\uparrow$
w/o TG-att 0.502 0.567 0.576 0.597 8.021
iso 0.499 0.553 0.568 0.583 8.788
noniso 0.489 0.547 0.566 0.581 9.483
ours 0.480 0.545 0.561 0.580 9.456
Table 4: Ablations on the AMASS dataset [52].
@@
We are also interested in evaluating how SkeletonDiffusion  generalizes to out-of-distribution motions. Hence, we test the methods trained on AMASS on unseen, real-life scenes from 3DPW[76] and report results in Tab. 3. We notice that, while CoMusion’s limb jitter between consecutive frames has worsened in the zero-shot setting, our method shows solid results and consistently the best body realism. We report a qualitative example in Fig. 6. CoMusion’s predictions appear diverse but present low semantic consistency with the input past. Furthermore, both predictions are humanly feasible as they present limb twisting or excessive bending.
@@
In Tab. 4, we ablate the main components of SkeletonDiffusion on AMASS. We compare the effect of TG-Attention layers on isotropic diffusion ($\boldsymbol{\Sigma}_{N}=I$ and ${\gamma}_{t}=0$) and analyze nonisotropic diffusion with our covariance reflecting joint connections $\boldsymbol{\Sigma}_{N}$ (Eq. 4) in the variant where ${\gamma}_{t}=1$ (as in Eq. 3) and our blending with the scheduler ${\gamma}_{t}$ (Eq. 5). We see that our TG-Attention layers already improve the TG-CNN architecture in the conventional isotropic diffusion paradigm. While the simple nonisotropic variant achieves state-of-the-art performance (see Tab. 1), our formulation with the scheduler ${\gamma}_{t}$ further improves the metrics and in particular, precision. Ablation results regarding the choice of connectivity matrix for $\boldsymbol{\Sigma}_{N}$ and its normalization are reported in the Sup.Mat., where we also show that our nonisotropic formulation requires fewer parameters and training epochs than the isotropic one.

### v2

@@ 4.3 SkeletonDiffusion
SkeletonDiffusion implements our nonisotropic diffusion formulation (Sec. 4.1) in latent space (Sec. 4.2). To obtain an explicit prior on realistic motions, we embed the knowledge about skeletal connections also in the network architecture.
@@
Applying our nonisotropic diffusion in latent space requires retaining the semantic meaning of each body joint (Sec. 4.2). We choose a fully GCN architecture and perform graph attention on the skeleton joints via Typed-Graph Convolutions [67]. For each layer taking as input features ${\boldsymbol{x}}\in\mathbb{R}^{J\times D_{in}}$, we define a feature extraction matrix $\mathbf{W}^{j}\in\mathbb{R}^{D_{in}\times D_{out}}$ for each joint $j$ with shared weights depending on the specific joint, and a feature aggregation matrix $\mathbf{G}\in\mathbb{R}^{J\times J}$. The features $\mathbf{f}\in\mathbb{R}^{J\times D_{out}}$ are first extracted for each joint $j$ independently as $\hat{\mathbf{f}}$ and then aggregated through
$\mathbf{f}=\mathbf{G}\cdot\hat{\mathbf{f}},\,\text{with}\,\,\hat{\mathbf{f}}^{j}=\mathbf{W}^{j}\cdot{\boldsymbol{x}}^{j}.$ (9)
We further define multi-head self-attention [75] on a joint level as Typed-Graph Attention and chose it as the architecture of the denoiser network. Both encoder and decoder are GRUs, exploiting the convenient inductive biases of recurrent neural networks for motion modeling [50] (Appendix B). With such architecture, the prior on the body joints is explicitly encoded in every layer.
Autoencoder and Latent Space Training
Given an input motion $\textbf{M}=\textbf{Y}_{0:\tilde{F}}\in\mathbb{R}^{\tilde{F}\times J\times 3}$ of arbitrary length $\tilde{F}\sim{\mathcal{U}}\{1,F\}$, the encoder $\mathrm{{e}}$ compresses the complex temporal information into latent space variables $\boldsymbol{z}=\mathrm{{e}}(\textbf{M})\in\mathbb{R}^{J\times L}$, where the joint dimension $J$ is kept intact and the latent dimension $L$ contains both temporal and spatial information. The decoder $\mathrm{{d}}$ learns to reconstruct the latent variable into a motion $\tilde{\textbf{M}}=\mathrm{{d}}(\boldsymbol{z},\textbf{X}_{-2:0})$. Here, conditioning the decoder on the previous two frames encourages smooth transitions between past and future [5]. The autoencoder is trained to reconstruct a motion according to the objective:
$L_{\mathrm{autoenc}}=L_{\mathrm{rec}}(\mathrm{{e}}(\textbf{M}),\textbf{M},\textbf{X}_{-2:0}),$ (10)
@@
$L_{\mathrm{rec}}(\boldsymbol{z},\textbf{M},\textbf{X}_{-2:0}):=\|\mathrm{{d}}(\boldsymbol{z},\textbf{X}_{-2:0})-\textbf{M}\|_{1}.$ (11)
We aim for a strong temporal representation, and let the latent space learn a general motion distribution of arbitrary length, fitting both observation and future motions. To avoid collapse towards the motion mean of the training data [9, 80], we employ curricular learning [8, 80, 1].
Figure 4: Qualitative Results on AMASS [53]. On the top, we report the input past observation and the ground truth future. The following rows display the corresponding predictions for each method: on the right, the closest to GT according to ADE, and on the two left–most columns, the two furthermost. Our closest competitors generate realistic motions but do not include a motion close to the GT (BeLFusion) or present evident unrealistic artifacts (CoMusion). Our method is the only one to produce realistic and diverse motions.
Latent Nonisotropic Diffusion
In latent space, the denoising network $\mathrm{{g}}$ learns via our nonisotropic diffusion formulation to denoise true latent variables $\boldsymbol{z}$ conditioned on past observations $\mathrm{{e}}(\textbf{X})$. Instead of predicting the noise ${\boldsymbol{\epsilon}}_{t}$ [31, 64], we directly approximate the true latent code ${\boldsymbol{x}}_{0}:=\boldsymbol{z}$ [63, 5] as ${\boldsymbol{x}}_{\boldsymbol{\theta}}=\mathrm{{g}}({\boldsymbol{x}}_{t},\mathrm{{e}}(\textbf{X}),t)$. To implicitly enforce diversity [5, 69, 29], we relax Eq. (8) by sampling $k=50$ predictions at each iteration and backpropagating the gradient only through the sample closest to the GT:
@@
Instead of choosing the sample that minimizes the diffusion loss [5, 69, 29], we choose the prediction that minimizes the reconstruction loss Eq. 11, finding that this benefits diversity in the generated ensemble. At inference, we denoise multiple latent codes ${\boldsymbol{x}}_{\theta}$ according to our reverse formulation LABEL:eq:reverse_diff. The generated latent codes are then decoded into future predictions $\tilde{\textbf{Y}}$.
@@ 5.1 Experimental Settings
We evaluate on the large-scale dataset AMASS [53] according to the cross-dataset evaluation protocol [5, 69]. We aim to test SHMP methods with real-world data obtained not from MoCap but from noise sources (e.g., RGB cameras, and sparse IMUs). To this end, we perform zero-shot experiments on 3D Poses in the Wild (3DPW) [78] for models trained on AMASS, and adapt the recent in-the-wild, large-scale dataset FreeMan[81] to the motion prediction task and retrain on it various state-of-the-art methods. We deem the conventionally employed Human3.6M dataset [34] less representative (only 7 subjects) and discuss it directly in Sec. F.2. As in previous works, we predict the next 2s into the future from observations of 0.5s.
@@
Recent SHMP works concentrate on four factors: precision, coverage of the ground truth test distribution (multimodal metrics), diversity, and realism. We employ conventional metrics [91, 5] and report their definition in Sec. D.1. While the CMD metric addresses realism, it is solely expressed in terms of joint velocities. The actual Body realism, e.g., bone lengths preservation along the motion, although crucial for meaningful predictions, is overlooked. Even worse, artifacts such as changes in limb lengths over time (limb stretching) and frequent inconsistencies between consecutive frames (limb jitter) impact other metrics, for example, by causing more diversity in the predictions, and so higher APD value (further experiments in Sec. F.1). This motivates us to investigate this aspect and propose new metrics. Given a future ground truth sequence Y with $\mathrm{B}$ limbs (or bones) and a predicted sequence $\tilde{\textbf{Y}}$, for each frame $\tau$ of the prediction associated pose $\tilde{\textbf{p}}_{\tau}$, we denote the length of the $j$-th limb as $\tilde{b}^{j}_{\tau}\in\mathbb{R}$. With $b^{j}\in\mathbb{R}$ being the ground truth length of the $j$-th limb, we define the normalized $j$-th limb length error $e^{j}_{\tau}$ and limb jitter $v^{j}_{\tau}$ at a time $\tau$ as:
$e^{j}_{\tau}:=\frac{1}{b^{j}}\left|b^{j}-\tilde{b}^{j}_{\tau}\right|,\quad v^{j}_{\tau}:=\frac{1}{b^{j}}\left|\tilde{b}^{j}_{\tau+1}-\tilde{b}^{j}_{\tau}\right|.$ (13)
By calculating the mean and root mean square error (RMSE) of $e^{j}_{\tau}$ and $v^{j}_{\tau}$ over the time dimension, we define four body realism metrics: mean for stretching str and jitter jit, and analogously RMSE. We also introduce the mean angle error (MAE) as complementary precision metric.
@@
ADE $\downarrow$ FDE $\downarrow$ MAE $\downarrow$ MMADE $\downarrow$ MMFDE $\downarrow$ APD $\uparrow$ CMD $\downarrow$ str jit str jit
Zero-Velocity 0.603 0.835 9.841 0.687 0.865 0.000 14.734 6.37 0.00 6.37 0.00
HumanMAC [16] 0.415 0.511 8.630 0.537 0.600 5.426 2.025 7.91 1.49 11.89 1.84
BeLFusion [5] 0.420 0.495 8.494 0.496 0.516 5.209 6.306 10.46 0.41 11.93 0.54
CoMusion [69] 0.389 0.480 7.812 0.527 0.525 6.687 2.764 7.94 0.81 10.27 1.05
SkeletonDiffusion (Ours) 0.374 0.457 7.424 0.506 0.508 6.732 3.166 7.58 0.51 9.64 0.66
@@ 5.2 Results
Following the cross-dataset evaluation protocol [5], we train on a subset of datasets belonging to AMASS and test on others (Tab. 1). Starting from the conventional metrics evaluation, our method already achieves state-of-the-art performance on the majority of the metrics, with a significant improvement on precision. Among other Diffusion-based methods (DM), SkeletonDiffusion and CoMusion contend with each other for first and second place according to diversity, realism, and multimodal metrics. Interestingly, the MAE values for DLow and GSPS do not reflect the performance ranking of the other precision metrics, while instead this holds for DM methods. Although VAE-methods tend to show higher diversity values such as of [84, 91], as already mentioned by previous works [5, 16], these values may often be the consequence of unrealistic motions with irregularities between past and future or inconsistent speed. From the qualitative example reported in Fig. 4, we notice that both the most diverse predictions of DivSamp represent a cartwheeling motion. While such motions may geometrically be diverse from each other and thus increase the APD, they are not only not semantically diverse but also not consistent with the past observation. Instead, our predictions are diverse at no expenses of realism (see also Fig. 1).
@@
Figure 7: Motions’ validity on different error tolerance on AMASS [53]. For every method, we show the evolution of valid motions quantity (y-axis) for which the maximal error is below a given threshold (x-axis). SkeletonDiffusion presents consistently the highest number of valid poses. CoMusion and VAE methods cannot generate predictions with an error lower than 2.5%.
@@
ADE $\downarrow$ FDE $\downarrow$ MAE $\downarrow$ MMADE $\downarrow$ MMFDE $\downarrow$ APD $\uparrow$ CMD $\downarrow$ str jit str jit
Alg Zero-Velocity 0.755 1.011 7.294 0.777 1.013 0.000 40.695 0.00 0.00 0.00 0.00
VAE TPK [79] 0.648 0.701 9.963 0.665 0.702 9.582 13.136 7.61 0.36 10.02 0.51
DLow [91] 0.581 0.649 8.820 0.602 0.651 13.772 11.977 8.53 0.42 11.28 0.61
GSPS [55] 0.552 0.650 8.469 0.578 0.653 11.809 12.722 6.38 0.29 8.65 0.35
DivSampling [19] 0.554 0.678 7.647 0.593 0.686 24.153 46.431 11.04 0.78 16.31 1.01
DM BeLFusion [5] 0.493 0.590 6.727 0.531 0.599 7.740 17.725 6.47 0.22 7.96 0.29
CoMusion 0.477 0.570 6.830 0.540 0.587 11.404 7.093 4.01 0.38 5.54 0.50
DM SkeletonDiffusion (Ours) 0.472 0.575 6.025 0.535 0.594 9.814 10.474 3.02 0.17 4.16 0.23
@@
We are also interested in evaluating how SkeletonDiffusion  generalizes to out-of-distribution motions. Hence, we test the methods trained on AMASS on unseen, real-life scenes from 3DPW [78] and report results in Tab. 3. We notice that, while CoMusion’s limb jitter between consecutive frames has worsened in the zero-shot setting, our method shows solid results and consistently the best body realism. We report a qualitative example in Fig. 6. CoMusion’s predictions appear diverse but present low semantic consistency with the input past. Furthermore, both predictions are humanly unfeasible as they present limb twisting or excessive bending.
Long Term Prediction and Challenging Scenario
We autoregressively feed generated motions to obtain a forecasting of 5s out of models trained to predict 2s (Appendix Tab. 11). We also design a challenging scenario testing on zero-shot generalization and noisy input data simultaneously (Appendix Tab. 10). In both settings we maintain the best realism with a significant gap and showcase state-of-the-art precision and diversity: the explicit inductive bias of SkeletonDiffusion  on the human body structure allows our method to preserve the body realism over time and generalize robustingly to noise and actions unseen at train time.
@@
In the Appendix Tab. 7, we report the ablations for the main components of SkeletonDiffusion on AMASS. Our TG-Attention layers improve the GCN architecture in the conventional isotropic diffusion paradigm. While the simple nonisotropic variant of Eq. 3 achieves state-of-the-art performance, our formulation with the scheduler ${\gamma}_{t}$ further improves the metrics and in particular, precision. Ablation results regarding the choice of connectivity matrix for $\boldsymbol{\Sigma}_{N}$ and its normalization are reported in Sec. E.2. We also show (Sec. E.3) that our nonisotropic formulation requires fewer parameters and training epochs than the isotropic one.